Find a.b given that A= 3m and B = 5m at an angle of 45°

1
Expert's answer
2022-01-07T09:29:19-0500

By definition, the dot product is given as follows:


"\mathbf{a}\cdot\mathbf{b} = AB\cos \theta"

where "\theta = 45\degree". Thus, obtain:


"\mathbf{a}\cdot\mathbf{b} = 3m\cdot 5m\cdot \cos 45\degree = 15m^2\cdot \dfrac{\sqrt{2}}{2} = 7.5\sqrt2m^2\approx 11m^2"

Answer. "7.5\sqrt2m^2\approx 11m^2".